Rule 1: Simple Probability
If a trial has k equally likely outcomes,
Give this one a try later!
the probability of any individual outcome is 1/k.
Example: What is the probability of rolling a 2 on a fair, 6-sided die?
Answer: 1/6
At CSULB, do male and female students show different levels of school spirit? I've
often wondered
about this when I walk across campus. On one of my walks I noticed that 7 out of 100
randomlysampled
students were male and were wearing CSULB attire. On another one of my walks I
noticed that
1 in 10 female students were wearing CSULB attire. It is also known that the student
,body is 57%
female. Based on these data, answer questions 27-30.
20) What is the probability that a randomly selected student is male and not wearing
any CSULB attire?
Give this one a try later!
0.36
2) The mode for a set of heights of students (grouped into 4cm classes) is determined.
What exactly is
the mode?
Give this one a try later!
The height that occurs most frequently
_________ are characteristics that differ among individuals
Give this one a try later!
variable
How do you expect each of these to change as sample size increases?
Give this one a try later!
, Mean same
variance same
sd same
standard error should dec
confidence interval should dec
Inclusion Interval Width same
16) In a certain normally distributed population with mean equal 15cm, 80% of the
central area under the curve lies in the range of 10 to 20 cm. What is the probability of
randomly selecting a variate (=observation measured value) measuring less than 10
cm?
Give this one a try later!
0.10
+kurtosis
Give this one a try later!
= leptokurtosis = data have higher central peak (higher concentration of
observations)
Location Statistics
Give this one a try later!
, tells you where on number line the sample values are located
includes: mean, mode, median, quartiles
Rule 2: Total Probability
Sum of the probabilities of all, mutually exclusive, possible outcomes
Give this one a try later!
equals 1
Example: the total probability of rolling 1 through 6 on a fair, 6-sided die is:
1/6 + 1/6 + 1/6+ 1/6 + 1/6 + 1/6 = 1.0
As the number of events increases for a given mean,
Give this one a try later!
the probability rapidly approaches zero
CHAPTER 7 - POISSON
Give this one a try later!
individuals/events occur randomly over space/time
Calculate the absolute expected frequency by
If a trial has k equally likely outcomes,
Give this one a try later!
the probability of any individual outcome is 1/k.
Example: What is the probability of rolling a 2 on a fair, 6-sided die?
Answer: 1/6
At CSULB, do male and female students show different levels of school spirit? I've
often wondered
about this when I walk across campus. On one of my walks I noticed that 7 out of 100
randomlysampled
students were male and were wearing CSULB attire. On another one of my walks I
noticed that
1 in 10 female students were wearing CSULB attire. It is also known that the student
,body is 57%
female. Based on these data, answer questions 27-30.
20) What is the probability that a randomly selected student is male and not wearing
any CSULB attire?
Give this one a try later!
0.36
2) The mode for a set of heights of students (grouped into 4cm classes) is determined.
What exactly is
the mode?
Give this one a try later!
The height that occurs most frequently
_________ are characteristics that differ among individuals
Give this one a try later!
variable
How do you expect each of these to change as sample size increases?
Give this one a try later!
, Mean same
variance same
sd same
standard error should dec
confidence interval should dec
Inclusion Interval Width same
16) In a certain normally distributed population with mean equal 15cm, 80% of the
central area under the curve lies in the range of 10 to 20 cm. What is the probability of
randomly selecting a variate (=observation measured value) measuring less than 10
cm?
Give this one a try later!
0.10
+kurtosis
Give this one a try later!
= leptokurtosis = data have higher central peak (higher concentration of
observations)
Location Statistics
Give this one a try later!
, tells you where on number line the sample values are located
includes: mean, mode, median, quartiles
Rule 2: Total Probability
Sum of the probabilities of all, mutually exclusive, possible outcomes
Give this one a try later!
equals 1
Example: the total probability of rolling 1 through 6 on a fair, 6-sided die is:
1/6 + 1/6 + 1/6+ 1/6 + 1/6 + 1/6 = 1.0
As the number of events increases for a given mean,
Give this one a try later!
the probability rapidly approaches zero
CHAPTER 7 - POISSON
Give this one a try later!
individuals/events occur randomly over space/time
Calculate the absolute expected frequency by